OpenStudy (anonymous):

how do I solve -4x^2+5x=1

OpenStudy (anonymous):

i would start with \[4x^2-5x+1=0\] and then try to factor

OpenStudy (anonymous):

since the only way to get two numbers whose product is 1 is \(1\times 1\) or \(-1\times -1\) i would try \[(4x-1)(x-1)=0\]

OpenStudy (anonymous):

thank you so much